A = 5.7 B = 4.3 4. Find the magnitude of vector C = A - B. Use all the information presented by the graph. A. 5.7 B. 6.9 C.7. the correct answer has been chosen, i just need the step by step solution...thanks
Answer #1

A -5.7 x B=403 C = A-B . . A = 50+ sin 38 1 +: 5.7 Cas 38°j A: 3.509 i + 4.4911 B = 4.3 Sin 24 i -3 C0924 B = 11749.- 3.93;